calculate the equilibrium level of income if C=R100 million+0.8Y and I=R125 million.
1
Expert's answer
2019-08-16T09:14:26-0400

Y = C + I + G

Y =R100 million+0.8Y + R125 million.

Y= 100+0.8Y+125=225+0.8Y

0.2Y=225

Y= R1125 million